OpenStudy (anonymous):

Let f(x) = \frac{1}{\sqrt{ x - 5 }} ,

OpenStudy (anonymous):

input to squares must be non negative for real solution so x can not be less than. 5

OpenStudy (anonymous):

do the work yourself lazy feather

OpenStudy (anonymous):

b*tch

OpenStudy (anonymous):

\[f(x)={1\over\sqrt{x-5}}\]what about it? :)

OpenStudy (austinl):

\(\large{f(x) = \dfrac{1}{\sqrt{ x - 5 }}}\) That is a function, what are we supposed to do with this?

OpenStudy (anonymous):

@brooklyn1, do you still need help with this?

OpenStudy (anonymous):

I'm trying to find the domain and range of f(x)

OpenStudy (anonymous):

I'm not sure where to begin

OpenStudy (anonymous):

It says I have to find both domain and range algebraically

OpenStudy (anonymous):

okay well when finding the domain, we can see that x cannot be less than or equal to 5, yes? because the radical is in the denominator, if x is 5, we will get a zero and that would make the function undefined. it can't be less than 5 because that would make the radical negative and that would be imaginary so we know domain begins after 5 \(\implies~domain=(5,\infty)\) now since the degree of the numerator is less than the denominator, we follow the asymptote rule and get that there is an asymptote at x=0 so there is a break in the range at 0 \(\implies~range=(-\infty,0)\cup(0,\infty)\)

OpenStudy (anonymous):

Thank you so much yummydum

OpenStudy (anonymous):

No problem!